OpenStudy (anonymous):

what is 10/12 + 4/9 =

OpenStudy (anonymous):

hi suzi. how old are u?

OpenStudy (anonymous):

Hi Anfar Im 29 and trying to help my son with his homework

OpenStudy (anonymous):

ok. first thing u should do is to make both fractions have the same denominator.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

u got how to finish it?

OpenStudy (anonymous):

one person got 2/3 and 1 and 5/18 we are confused who is right?

OpenStudy (anonymous):

i think it should be 23/18

OpenStudy (anonymous):

reduced is 1 5/18 right ?

OpenStudy (anonymous):

to make the denominator the same, note that: 10/12=30/36 and 4/9=16/36 so now u have 30/36+16/36=(30+16)/36=46/36=23/18

OpenStudy (anonymous):

yea 23/18

OpenStudy (anonymous):

or 1/5/18

OpenStudy (anonymous):

one-fifth over eighteen

OpenStudy (anonymous):

sory, misread the 2nd answer. 1 5/18 is correct as well

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

anfar u a tutor

OpenStudy (anonymous):

im a 11th grade student, wit a 3.85gpa

OpenStudy (anonymous):

suzi: u r welcome natural: lol.no im not. only help what i can

OpenStudy (anonymous):

Wow thats really awesome (that you can help) math is not my strong point so thanks again

OpenStudy (anonymous):

no problem suzi. glad to help im taking math as my major. :)